OpenStudy (anonymous):

I have been trying to work this problem, it doesnt seem to have enough information. Chlorine is frequently used to disinfect swimming pools. The chlorine concentration should remain between 1.5 and 2.5 parts per million (ppm) for safe swimming. After a warm, sunny day only 80% of the chlorine may remain in the water, with the other 20% dissipating into the air or combining with chemicals in the water. Let model the concentration of chlorine in parts per million after t days. What is the initial concentration of chlorine in the pool?

OpenStudy (amistre64):

your equation got lost in translation...

OpenStudy (anonymous):

I m just tryin to figure out how much initial concentration of chlorine in the pool

OpenStudy (amistre64):

"Let (......) model the concentration of chlorine in parts" the equation is missing. When you copy and pasted your question it ate the equation because it didnt recognize it

OpenStudy (amistre64):

if it aint an equation; then it ate whatever was witting there to begin with

OpenStudy (anonymous):

f(t)=3(0.8)^t

OpenStudy (amistre64):

well, it looks like we revert this to a table of times and try to make a recursion equation out of it

OpenStudy (amistre64):

A{n+1} = A{n} (.8) +C

OpenStudy (amistre64):

t=1; A = 2.4 t=2; A = 1.92 t=3; A = 1.536 1.92 = 2.4(.8) + C perhaps?

OpenStudy (amistre64):

C = 1.92 - (2.4*.8) C = 1.92 - 1.92 C = 0 maybe that works out lol

OpenStudy (amistre64):

C = 1.536 - (.8*1.92) C = 1.536 - 1.536 ... seems to be good

OpenStudy (amistre64):

2.4 = A{0}*.8 2.4\.8 = A{0} A{0} = 3 the inital amount was at 3 parts per mill if i read it right

OpenStudy (anonymous):

Ok so I have another quetions mathematically could the concentration ever be zero

OpenStudy (amistre64):

no; since the inital equation you gave can never have a zero component; the answer can never be 0; 3(0) = 0 but; .8^t never equals 0 0(.8^t) = 0 but; 3 never equals 0 so conclusion is: it can never reach zero